Two point charges ares separated by a distance of one meter .The two point charges have magnitudes of 2.5x10^-4 C and 1.5 x10^-2 C.Find the electric force between the two. Round your final answer to it's nearest hundreths.

1
Expert's answer
2022-03-07T12:21:43-0500

Answer

Force between them

"F=\frac{9*10^9*2.5*10^{-4}*1.5* 10^{-2} }{1}\\=33.75*10^4N"